### v1

# VGC: A High-Performance Zone-Based Garbage Collector Architecture for Python with Partitioning and Parallel Execution

Abstract
The Virtual Garbage Collector (VGC) introduces a novel memory management framework designed to optimize performance across diverse systems, ranging from resource-constrained embedded devices to high-performance parallel architectures. Unlike conventional garbage collectors, VGC employs a dual-layer architecture—Active VGC and Passive VGC—to enable efficient, low-overhead memory management. Active VGC dynamically manages runtime objects using a concurrent mark-and-sweep strategy tailored for parallel workloads, reducing pause times by up to 30% compared to generational collectors in multithreaded benchmarks. Passive VGC, which operates at compile time, optimizes static object allocation through predictive memory mapping, minimizing fragmentation by aligning objects to cache boundaries. This separation of responsibilities ensures predictable memory access patterns, reduces total memory usage by up to 25%, and improves scalability for modern parallel applications. By integrating compile-time and run-time optimizations, VGC offers a robust, adaptable solution for memory-intensive systems, achieving superior performance and reliability in both low-level and high-level programming environments.
1 Introduction
Garbage collection is a fundamental component of modern programming languages, automating memory management to prevent resource leaks and enhance software reliability. In Python, however, the garbage collector’s reliance on reference counting and the Global Interpreter Lock (GIL)[3], introduces several performance bottlenecks, including increased memory overhead, excessive fragmentation, unpredictable collection pauses, and limited support for true parallel execution. These constraints significantly reduce Python’s scalability in high-performance computing scenarios, particularly in data-intensive and multithreaded workloads.Classical garbage collection techniques rely on heap traversal, reference counting, or generational heuristics, which introduce non-deterministic pause behavior and scaling limitations [5], To address these challenges, we propose the Virtual Garbage Collector (VGC), a novel memory management framework designed to augment or serve as an alternative to Python’s GIL-based garbage collector. VGC employs a dual-layer architecture consisting of Active VGC, which manages dynamic objects at runtime, and Passive VGC, which optimizes the allocation of static objects at compile time. This separation enables fine-grained and efficient memory handling by classifying objects into three distinct zones—Red, Green, and Blue—based on usage frequency and computational complexity. Frequently accessed objects are prioritized in the Green Zone, while rarely used or high-complexity objects are allocated to the Blue or Red Zones, reducing memory fragmentation and improving cache locality. VGC introduces a checkpoint-based bitfield tracking system that replaces per-object reference counting, allowing object states to be evaluated using bitwise logic operations (e.g., AND, OR, NOT) in constant time. This eliminates recursive traversal during garbage collection and reduces latency by up to 20% in initial benchmarks. Furthermore, a lightweight Yield Memory layer handles small-scale operations, such as primitive variable assignments and basic arithmetic, without triggering full garbage collection cycles to further minimize overhead. Unlike conventional garbage collectors (e.g., Java’s CMS or Go’s concurrent GC), VGC’s triadic zoning and checkpoint-based design provide a predictable, low-latency, and scalable memory management solution suitable for parallel workloads. This paper presents the architecture and operational principles of VGC, its RGB zoning strategy, and benchmark results demonstrating its advantages over Python’s GIL-based garbage collector. By integrating partitioned memory zones, bitfield tracking, and yield-based optimization, VGC achieves reduced memory usage, lower fragmentation, and enhanced scalability, paving the way for efficient memory management in modern computing environments.
2 System Architecture
The architecture of the Virtual Garbage Collector (VGC) is designed around a layered, zone-oriented framework that emphasizes efficiency, determinism, and parallel scalability. Unlike traditional garbage collectors, which depend on recursive traversal and per-object reference counting, VGC employs a compact bit addressing and bitfield checkpoint system that allows constant-time object state evaluation. This design enables VGC to act both as an alternative and as a parallel companion subsystem to the default Python memory manager.
2.1 Overview of Architectural Layers
VGC consists of four key layers, each responsible for a distinct phase of memory lifecycle management:
  1. 1.
    Memory Zoning Layer (R/G/B): All allocated objects are stored in one of three predefined memory zones—Red, Green, or Blue. This classification is based on access frequency, object complexity, and lifetime. Each zone functions as a semi-independent allocator with its own internal management structures, minimizing interference between workloads of different types.
  2. 2.
    Checkpoint Bitfield Layer: Instead of storing reference counts in each object header, VGC maintains a centralized checkpoint table in which each object is represented by a single bit or a small fixed bit pattern. These bits encode object states such as active, idle, recyclable, or scheduled for collection. State evaluation and updates are performed through logic gate operations (AND, OR, NOT, XOR, XNOR, NOR, NAND), enabling constant-time liveness detection.
  3. 3.
    Yield Memory Layer: This lightweight execution layer acts as a short-term buffer or workspace for small-scale operations such as variable assignments, simple arithmetic, and temporary objects. Yield memory can operate independently of the checkpoint and zoning layers for ephemeral workloads, reducing unnecessary GC overhead. Once execution completes, yield memory either discards the object or notifies the checkpoint system to register persistent values.
  4. 4.
    Control and Execution Layer: This layer orchestrates the interaction between zones, checkpoint logic, and yield memory. It includes scheduling logic, partitioning control for multi-core workloads, and optional active-passive GC layering. Active VGC manages runtime and dynamic workloads, whereas Passive VGC optimizes static allocations at compile time.
3 Working Principle
3.1 R/G/B Zone Model
VGC classifies each allocated object into one of three memory zones based on its observed access frequency $a(o)$ and mutation rate $\mu(o)$, evaluated through the checkpoint bitfield state.Unlike traditional generational garbage collectors that rely on object promotion across generations [2], VGC enforces a static-zone lifecycle in which objects expire and are reallocated rather than migrated.
  • Red Zone (R): Rarely-Accessed Objects.
    Contains objects with low access frequency and low mutation activity. These objects contribute minimally to the active working set and are reclaimed aggressively to avoid unnecessary memory retention.
  • Green Zone (G): Frequently-Accessed Objects.
    Represents the program’s hot working set. Objects in this zone exhibit high access frequency, low lookup latency, and are prioritized for fast allocation and direct access. This zone is optimized for predictable performance.
  • Blue Zone (B): Medium-Access Objects.
    Contains objects with intermediate access or mutation rates. These objects remain stable across execution epochs but are not part of the primary active working set. Blue-zone memory is compacted periodically and validated through sparse checkpoints.
Each zone maintains its own bit-address table, allowing constant-time object indexing without pointer traversal. This eliminates hierarchical heap walking, reduces cache misses, and enables deterministic memory access across parallel execution units. Objects do not migrate between zones; instead, expired objects are reclaimed and reallocated into their new target zone when required.
3.2 Yield Memory Execution Path
For ephemeral operations such as:
x = 85 // a random variable
the standard garbage collection pipeline is bypassed. Yield Memory allocates a minimal execution unit, executes the operation, and immediately releases the allocation.
By skipping both checkpoint evaluation and VGC zone allocation for short-lived workloads, Yield Memory provides a fast execution path that reduces allocation overhead and runtime latency.
If persistence is required, Yield Memory notifies the checkpoint system, which assigns a 3-bit liveness state and registers the object into the appropriate VGC zone (typically Green).
3.3 Three-Bit Checkpoint Architecture
The checkpoint layer forms the core of VGC’s memory tracking system. Unlike conventional reference counting, which maintains an integer per object, VGC uses a compact 3-bit encoding for each object’s state. This allows for richer state representation while preserving constant-time evaluation. This 3-bit architecture allows VGC to precisely track object lifecycle transitions without per-object counters. State changes are performed via bitwise logic operations rather than recursive reference updates.
S.No Bit State Meaning Action
1 000 Idle (not accessed recently) Wait / Sleep
2 001 Active (currently in use) Keep Alive
3 010 Candidate for promotion
(higher priority zone)
Evaluate
4 011 Candidate for demotion
(lower priority zone)
Evaluate
5 100 Persistent (long-lived object) Keep Stay
6 101 Deferred (lazy checkpoint,
waiting confirmation)
Defer Sweep
7 110 Marked (garbage collection
scheduled)
Prepare for
Deletion
8 111 Expired (no longer needed) Reclaim Immediately
Table 1: Three-Bit Checkpoint State Table showing corresponding meanings and actions.
3.4 Object Feature Profiling and Runtime Metrics
To determine the appropriate zone for each allocated object, VGC evaluates a set of measurable runtime features. These features reflect how the object behaves during execution and are updated through lightweight counters maintained by the interpreter and core threads.
For an object $o$, the following metrics are recorded:
  • $\lambda(o)$Allocation Rate (objects/second): The instantaneous rate at which objects of the same type are allocated during execution. This is reflected in the short-term intensity of object creation.
  • $\tau(o)$Lifetime Duration (seconds): The elapsed time between the object’s allocation and its last observed reachability checkpoint.
  • $\mu(o)$Mutation Rate (writes/second): The frequency of state updates to the object, indicating how actively the data are being modified.
  • a(o)a(o)Access Rate (reads/second): The number of read operations referencing the object over time, reflecting how frequently it is used.
  • s(o)s(o)Object Size (bytes): The total allocated memory footprint of the object.
  • $\rho(o)$Referential Fan-Out: The number of distinct inbound references point to the object, representing its connectivity within the program’s object graph.
  • $\chi(o)$Computational Complexity Weight: A dimensionless factor representing the expected cost of evaluating,in a check-point system, or releasing the object during garbage collection.
These metrics are tracked using exponential moving averages to ensure stability over time, allowing VGC to classify and manage objects deterministically without iterative heap scanning or recursive reference traversal.
3.5 Zone Allocation Function
  • Let $a(o)$ denote the access frequency of object $o$ over a checkpoint window.
  • Let $\mu(o)$ denote the mutation rate of $o$ over the same window.
  • Let $a_{R}<a_{G}$ be the access thresholds.
  • Let $\mu_{R}<\mu_{G}$ be the mutation thresholds.
  • Let $C_{z}(o)$ denote the cost model of placing $o$ in zone $z\in\{R,G,B\}$.
function of Zone Allocation:
$Z(o)=\begin{cases}R,&\text{if }a(o)<a_{R}\ \land\ \mu(o)<\mu_{R},\\[4.0pt] G,&\text{if }a(o)\geq a_{G}\ \lor\ \mu(o)\geq\mu_{G},\\[4.0pt] B,&\text{otherwise, with tie-breaker }\arg\min_{z}C_{z}(o).\end{cases}$
where $C_{z}(o)$ is the cost model of placing $o$ in zone $z$.
3.5.1 Boundary Cases
  • At $a(o)=a_{R}$ or $\mu(o)=\mu_{R}$, the condition for $R$ is not met (strict inequalities), so the object is not assigned to $R$.
  • At $a(o)=a_{G}$ or $\mu(o)=\mu_{G}$, the condition for $G$ is met (non-strict inequalities), so the object is assigned to $G$.
  • If $a_{R}\leq a(o)<a_{G}$ and $\mu_{R}\leq\mu(o)<\mu_{G}$, the object falls into the “otherwise” case and is assigned to $B$, except that any ambiguity is resolved by the arg min tie-breaker on $C_{z}(o)$:
    $Z(o)=\arg\min_{z\in\{R,G,B\}}C_{z}(o).$
3.6 Allocation Decision Function
To determine the appropriate memory zone for an object $o$, VGC evaluates its runtime features:
$\lambda(o),\ \tau(o),\ \mu(o),\ a(o),\ s(o),\ \rho(o),\ \chi(o)$
zone-selection thresholds:
$\Theta=\{\tau_{R},\tau_{G},\ \mu_{R},\mu_{G},\ a_{R},a_{G},\ s_{R},s_{G}\}$
with the ordering constraints:
$\tau_{R}<\tau_{G},\quad\mu_{R}>\mu_{G},\quad a_{R}>a_{G},\quad s_{R}<s_{G}.$
$Z(o)=\begin{cases}R,&\tau(o)\leq\tau_{R}\land\mu(o)\geq\mu_{R}\land a(o)\geq a_{R}\land s(o)\leq s_{R},\\[4.0pt] G,&\tau_{R}<\tau(o)\leq\tau_{G}\land\mu_{G}\leq\mu(o)<\mu_{R}\land a_{G}\leq a(o)<a_{R}\land s_{R}<s(o)\leq s_{G},\\[4.0pt] B,&\tau(o)>\tau_{G}\lor\mu(o)<\mu_{G}\lor a(o)<a_{G}.\end{cases}$
If multiple conditions apply, a tie-breaking rule selects the lowest expected cost zone:
$Z(o)=\arg\min_{z\in\{R,G,B\}}C_{z}(o),$
where $C_{z}(o)$ is the expected per-collection cost of storing the object $o$ in zone $z$ .
3.6.1 Runtime Features and Thresholds
For an object $o$, define its measured features:
$f(o)=\left(\lambda(o),\ \tau(o),\ \mu(o),\ a(o),\ s(o),\ \rho(o),\ \chi(o)\right).$
Let the zone-selection thresholds be
$\Theta=\{\tau_{R},\tau_{G},\ \mu_{R},\mu_{G},\ a_{R},a_{G},\ s_{R},s_{G}\}$
with ordering constraints:
$\tau_{R}<\tau_{G},\qquad\mu_{R}>\mu_{G},\qquad a_{R}>a_{G},\qquad s_{R}<s_{G}.$
These constraints enforce monotone separation among the zones.
3.6.2 Eligibility Predicates
Define the zone-eligibility predicates:
$E_{R}(o):\ \tau(o)\leq\tau_{R}\ \wedge\ \mu(o)\geq\mu_{R}\ \wedge\ a(o)\geq a_{R}\ \wedge\ s(o)\leq s_{R},$
$E_{G}(o):\ \tau_{R}<\tau(o)\leq\tau_{G}\ \wedge\ \mu_{G}\leq\mu(o)<\mu_{R}\ \wedge\ a_{G}\leq a(o)<a_{R}\ \wedge\ s_{R}<s(o)\leq s_{G},$
$E_{B}(o):\ \tau(o)>\tau_{G}\ \vee\ \mu(o)<\mu_{G}\ \vee\ a(o)<a_{G}\ \vee\ s(o)>s_{G}.$
Define the set of applicable zones:
$\mathcal{A}(o)=\{\,z\in\{R,G,B\}\mid E_{z}(o)\text{ is true}\,\}.$
3.7 Zone Cost Model
Let $C_{\text{mark}}(o)$ be the cost to mark object $o$, $C_{\text{scan}}(o)$ the cost to scan its references, and $C_{\text{stage}}(o)$ the cost associated with promotion or demotion (i.e., expiration and reallocation rather than in-place relocation).
The per-zone cost is parameterized as:
$C_{R}(o)=\underbrace{\alpha_{R}\chi(o)}_{\text{mark}}+\underbrace{\beta_{R}\rho(o)}_{\text{scan}}+\underbrace{\gamma_{R}s(o)}_{\text{stage (reclaim)}}$
$C_{G}(o)=\alpha_{G}\chi(o)+\beta_{G}\rho(o)+\gamma_{G}s(o)$
$C_{B}(o)=\alpha_{B}\chi(o)+\beta_{B}\rho(o)+\gamma_{B}s(o)$
with policy weights:
$\gamma_{R}>\gamma_{G}>\gamma_{B},\qquad\alpha_{R}\approx\alpha_{G}>\alpha_{B},\qquad\beta_{R}\geq\beta_{G}\geq\beta_{B},$
reflecting that:
  • Red-zone objects are short-lived and frequently reclaimed.
  • Green-zone objects form the stable working set with moderate updates.
  • Blue-zone objects are long-lived and rarely re-staged except during periodic compaction.
The expected pause contribution per collection for zone $z$ is:
$P_{z}=\sum_{o\in z}C_{z}(o)\cdot\pi_{z},$
where $\pi_{z}\in(0,1)$ controls the fraction of $C_{z}(o)$ executed during stop-the-world versus concurrent phases.
3.7.1 Definitions and Scope
  • Let $o$ denote an object.
  • Let $C_{\text{mark}}(o)$ be the cost to mark $o$, $C_{\text{scan}}(o)$ the cost to scan $o$’s references, and $C_{\text{stage}}(o)$ the cost associated with promotion/demotion (i.e., expiration and reallocation rather than in-place relocation).
  • The model per-zone costs using nonnegative parameters and additive feature functions. Units are arbitrary “work units” per collection; parameters absorb constant factors of the implementation.
$C_{z}(o)=\alpha_{z}\chi(o)+\beta_{z}\rho(o)+\gamma_{z}s(o)\quad\text{for }z\in\{R,G,B\}$
3.7.2 Feature functions
  • $\chi(o)$: mark-related work attributable to $o$ (e.g., number of header touches).
  • $\rho(o)$: scan-related work for $o$’s references (e.g., references traversed).
  • $s(o)$: staging work for $o$ (e.g., bytes or objects expired/reclaimed during (promotion/demotion).
All three are nonnegative and dimensioned in primitive “units” that the parameters convert into cost.
3.7.3 Per-Zone Cost Parameterization
For a zone $z\in\{R,G,B\}$, let the zone-specific parameters $\alpha_{z},\beta_{z},\gamma_{z}\geq 0$ control the weighting of the mark, scan, and stage components:
$C_{z}(o)=\alpha_{z}\,\chi(o)+\beta_{z}\,\rho(o)+\gamma_{z}\,s(o).$
  • The first term $\alpha_{z}\chi(o)$ captures mark complexity.
  • The second term $\beta_{z}\rho(o)$ captures scan overhead.
  • The third term $\gamma_{z}s(o)$ captures stage cost, corresponding to expiration and reallocation rather than in-place movement.
3.7.4 Estimation Notes
  • The parameters $\{\alpha_{z},\beta_{z},\gamma_{z}\}$ are typically learned through nonnegative least-squares fitting data from the per-collection profiling data.
  • Identifiability improves when features $\chi(o),\rho(o),s(o)]$ vary between objects and checkpoint intervals.
  • To maintain interpretability, feature magnitudes should be scaled to similar ranges, or regularization may be applied.
$C_{z}(o)=\alpha_{z}\chi(o)+\beta_{z}\rho(o)+\gamma_{z}s(o),\qquad z\in\{R,G,B\}.$
3.8 16-Byte Alignment and VGC Checkpoint Indexing
CPython aligns all heap objects to a 16-byte boundary due to its internal object layout.Each object begins with a fixed 16-byte header containing the reference count and a pointer to its type descriptor. Small-object allocations are rounded to the nearest 8 or 16 bytes by the pymalloc allocator, ensuring consistent alignment across the heap.
This alignment property is critical to VGC. Because objects occupy uniformly aligned memory regions, the Virtual Garbage Collector represents each object using a compact 3-bit entry in a global checkpoint bitfield. The index of an object in the checkpoint table can therefore be computed in constant time using integer bit shifts rather than pointer dereferencing or graph traversal.
This enables the VGC to perform object state evaluation using logic-gate operations without scanning the heap, providing predictable and deterministic garbage collection behavior.
3.8.1 Zone-Index Mapping and Memory Layout Specification
To operationalize the checkpoint-based evaluation model, the Virtual Garbage Collector (VGC) defines a deterministic mapping from physical object addresses to checkpoint bit indices and, subsequently, to the R/G/B memory zones. This mapping ensures constant-time zone lookup, predictable garbage-collection behavior, and correct parallel partition assignment.Given that all CPython-managed objects are aligned to 16-byte boundaries, each object’s index in the global checkpoint bitfield is computed as:
$\text{Index}(o)=\frac{\text{Address}(o)-\text{Base}}{16}.$ (1)
This index uniquely identifies the object and determines its placement within the zone-specific bitfield ranges defined below.
3.8.2 Global Zone Boundary Layout
Let:
$N_{R}$ $\ =\text{maximum entries in Red zone},$
$N_{G}$ $\ =\text{maximum entries in Green zone},$
$N_{B}$ $\ =\text{maximum entries in Blue zone}.$
VGC partitions the global checkpoint table into three contiguous regions:
$\text{CheckpointTable}=[R_{\text{bits}}\mid G_{\text{bits}}\mid B_{\text{bits}}].$
The index ranges for each zone are:
$\text{Red zone}:$ $\qquad i\in[0,\,N_{R}),$ (2)
$\text{Green zone}:$ $\qquad i\in[N_{R},\,N_{R}+N_{G}),$ (3)
$\text{Blue zone}:$ $\qquad i\in[N_{R}+N_{G},\,N_{R}+N_{G}+N_{B}).$ (4)
Thus, zone membership is determined in $O(1)$ time via boundary comparison. The zone decision formulas in Sections 3.5 and 3.6 determine which zone an object belongs to, while this layout determines where the object is placed.
3.8.3 VGC adopts a static-zone memory model
objects do not migrate between R/G/B zones after allocation. If checkpoint evaluation indicates that an object should belong to a different zone, the object undergoes:
  1. 1.
    expiration in its current zone.
  2. 2.
    reallocation into the new zone’s index range.
This ensures zone-boundary stability and avoids the complexity of in-place relocation.
3.8.4 Generational Subdivision Within Zones
Each zone is subdivided into three generational segments: Gen0, Gen1, and Gen2. For a zone $z$ of size $N_{z}$, the index ranges are:
$\text{Gen0}:$ $\qquad i\in[\text{start}_{z},\;\text{start}_{z}+\alpha N_{z}),$ (5)
$\text{Gen1}:$ $\qquad i\in[\text{start}_{z}+\alpha N_{z},\;\text{start}_{z}+\beta N_{z}),$ (6)
$\text{Gen2}:$ $\qquad i\in[\text{start}_{z}+\beta N_{z},\;\text{start}_{z}+N_{z}),$ (7)
where typical configuration values are $\alpha=0.25$ and $\beta=0.75$. These proportions are tunable based on workload characteristics.
3.8.5 Partition Mapping for Parallel Execution
For parallel execution, each zone $z$ is further divided into $P_{z}$ partitions. Partition $p$ is assigned a contiguous index interval:
$\text{Partition }p\text{ of zone }z:\quad i\in\left[\text{start}_{z}+\frac{p}{P_{z}}N_{z},\;\text{start}_{z}+\frac{(p+1)}{P_{z}}N_{z}\right).$ (8)
This design ensures lock-free operation, isolated memory access per core, and deterministic checkpoint evaluation. The 3-bit entries per object allow boundaries to align with machine-word sizes for SIMD-optimized bitfield processing.
3.8.6 Integration with Logic-Gate Checkpoint Evaluation
For any object $o$:
  1. 1.
    The address determines its index via Eq. (1).
  2. 2.
    The index determines its zone by boundary comparison.
  3. 3.
    The zone determines:
    • the bitfield region storing its 3-bit state,
    • the corresponding logic-gate evaluation chain (R, G, or B),
    • the partition—and thus the CPU core—responsible for its processing.
This establishes a complete mapping from object address to zone, generation, partition, and checkpoint evaluation path, all in constant time.
3.9 Active and Passive VGC Integration
VGC supports Active–Passive dual layering:
  • Active VGC — Handles dynamic runtime workloads such as recursion, loops, and asynchronous I/O.
  • Passive VGC — Handles compile-time or static allocations, aligning objects to stable memory regions.
This separation allows stable management of static objects while enabling rapid adaptation for runtime workloads. The checkpoint layer can differentiate between static and dynamic lifecycles through specific 3-bit states.
3.10 Parallelism and Partitioning
VGC integrates a partitioning model that distributes workloads across CPU cores, with zones operating independently. Checkpoint bitfield updates are parallelized, avoiding locking overhead typical in Python’s GIL. Because the checkpoint system is based on 3-bit state logic, partitioning can use bitmask operations to batch-evaluate objects across zones concurrently. This allows multiple threads or interpreters to process memory in parallel without race conditions or serialized refcount updates. In summary, VGC’s working principle combines three-bit checkpoint state encoding, zone-based allocation, logic-gate-driven decision making, lambda-optimized yield execution, and parallel Execution by partitioning workloads. This results in a deterministic, low-latency, and highly scalable alternative to Python’s GIL-based garbage collector.
4 Lambda and Three-Bit Checkpoint Mapping
Each lambda-evaluated object can be associated with a 3-bit state based on its persistence:
  • 000 – ephemeral lambda (immediate discard after evaluation).
  • 001 – active lambda (used in a running scope).
  • 100 – persistent lambda (checkpointed and stored).
  • 101 – deferred evaluation (lazy checkpoint, stored later).
Let:
  • $L$ denote the set of lambda operations,
  • $S$ the set of checkpoint states,
  • $Z$ the set of zones.
Then:
$\phi:L\to S\to Z$
The mapping function $\phi$ is defined as:
[…]
This mapping ensures that ephemeral tasks do not incur additional checkpoint or zone allocation costs, while persistent and deferred tasks are efficiently integrated into the VGC memory hierarchy
The use of lambda calculus in VGC therefore provides:
  • Minimal overhead for small, short-lived computations.
  • Deterministic behavior through explicit state mapping.
  • Seamless integration with the 3-bit checkpoint system.
5 Logic Gate–Based Architecture
The Logic Gate–Based Architecture[5], forms the computational core of the VGC checkpoint system. Rather than relying on iterative reference tracking or recursive object scanning, VGC encodes each object’s state within a compact 3-bit field, which is processed directly through gate-level logic operations. This design enables constant-time evaluation of object lifecycles and eliminates the overhead associated with conditional branching or traversal-based garbage collection.
5.1 Checkpoint Logic Integration
Each allocated object is represented by a 3-bit checkpoint code, denoted as:
$S=(s_{2},s_{1},s_{0})$
where each bit corresponds to a specific state flag used to determine object liveness and collection readiness. The checkpoint bitfields are grouped into logical blocks, allowing simultaneous evaluation of multiple objects using bitwise operations. Instead of maintaining numerical reference counts, the system interprets combinations of bits through logic gates that perform rapid evaluation at the binary level.
  • The AND gate ($A\land B$) ensures that an object remains valid only if both the current state and zone permission bits are active. It acts as a confirmation gate for “keep-alive” conditions.
  • The OR gate ($A\lor B$) preserves objects that are still linked to high-priority zones (such as Green or Blue) even when their primary reference becomes inactive.
  • The NOT gate ($\lnot A$) is used to invert inactive bits, identifying idle or expired objects ready for reclamation.
  • The XOR gate ($A\oplus B$) tracks state transitions by detecting temporal differences between consecutive epochs of the same object.
  • The XNOR gate ($\overline{A\oplus B}$) identifies stability across multiple evaluation cycles, ensuring that persistent objects are not prematurely collected.
  • The NAND and NOR gates are used in composite checkpoint circuits, enabling conditional sweeping and deferred reclamation for objects marked with state patterns $101$ (Deferred) or $110$ (Marked).
5.2 Gate-Level Evaluation Model
During every checkpoint cycle, the bitfield array is loaded into a low-level logic evaluation routine where each 3-bit pattern undergoes concurrent gate computation. The logic evaluation is expressed as:
$O=(S\land Z)\lor(\lnot S\land P)$
where:
  • $S$ represents the current state bits of the object.
  • $Z$ represents the zone activation mask ($R,G,B$).
  • $P$ represents pending or deferred state bits.
  • $O$ is the output that determines whether the object is still allocated or released.
This logical combination allows VGC to independently assess the readiness of each object without interacting with neighboring references or recursive memory links. Thus, multiple checkpoint evaluations can occur in parallel with no dependency chain, achieving $\mathcal{O}(1)$ time complexity per object.
5.3 Bitfield Processing and Memory Synchronization
All checkpoint bits are aligned to 16-byte memory boundaries due to the CPython design constraints to ensure cache-line efficiency and predictable access latency. The bitfield is divided into three synchronized regions corresponding to the R, G, and B zones:
$\text{Checkpoint Table}=[R_{\text{bits}}\,|\,G_{\text{bits}}\,|\,B_{\text{bits}}]$
Each region is independently evaluated through its own logic chain. For example:
$R^{\prime}$ $=(R\land\lnot B)\lor(R\land G)$
$G^{\prime}$ $=(G\lor R)\land\lnot B$
$B^{\prime}$ $=(B\land\lnot G)$
These relations prevent inter-zone conflicts and ensure stable memory boundaries across concurrent threads. The bitfield synchronization unit within VGC merges results through gate aggregation, maintaining coherence between Active and Passive VGC layers.
5.4 Advantages of Gate-Level Checkpoint Processing
  • Constant-Time Evaluation: Each logic computation occurs in a fixed number of operations, irrespective of heap size.
  • Parallel-Friendly: Bitfield arrays can be batch-evaluated across CPU cores without shared reference conflicts.
  • Deterministic Behavior: Gate-level execution produces consistent results under concurrent workloads.
  • Low Memory Overhead: Only 3 bits per object are required to encode all active states.
  • Hardware Alignment: Operations are naturally compatible with CPU bit arithmetic and SIMD instructions.
Through its logic gate–driven checkpoint evaluation, VGC transforms memory management into a design that bridges software-level garbage collection with hardware-inspired parallelism, creating a deterministic and high-performance foundation for Python’s next-generation runtime.
6 Partition Theory and Parallel Execution Model
The Partition Theory and Parallel Execution Model of the Virtual Garbage Collector (VGC) establishes a mathematical and architectural foundation for distributing workloads across multiple CPU cores and threads instances. This mechanism enables VGC to achieve deterministic, scalable parallelism while maintaining memory consistency across all active threads and cores. While scalable allocators such as Hoard reduce contention in multithreaded environments [1], they operate below the runtime layer and do not address object lifecycle semantics or interpreter-level execution parallelism.
6.1 Partition and Parallel Execution (PPE) Runtime Architecture
Partition and Parallel Execution (PPE) is the runtime execution architecture[4], responsible for decomposing workloads into independent partitions and executing them concurrently across available CPU cores and hardware threads. PPE operates independently of memory reclamation logic and serves as the execution backbone upon which the Virtual Garbage Collector (VGC) is layered.
The primary objective of PPE is to eliminate underutilization of processor resources by ensuring that no CPU core remains idle when parallelizable work is available. Unlike Python’s Global Interpreter Lock (GIL), which serializes bytecode execution, PPE enables true multi-core execution by design.
PPE targets computational patterns that are traditionally constrained by interpreter-level serialization, including:
  • Loop-intensive workloads
  • Recursive and deep-recursive execution
  • Matrix and numerical computation
  • Interpreter-driven task dispatch
Each workload is partitioned into logically independent execution units based on the processor’s core and thread topology. These partitions are scheduled for parallel execution using native OS threads with explicit core affinity, ensuring predictable performance scaling across dual-core, quad-core, and higher-core-count systems.
Formally, PPE maps the total workload $W$ into $n$ partitions such that:
$W=\bigcup_{i=1}^{n}W_{i},\quad W_{i}\cap W_{j}=\emptyset\text{ for }i\neq j$
where each partition $W_{i}$ is executed concurrently on a dedicated core or hardware thread.
PPE is deliberately lightweight and avoids global synchronization barriers. Inter-partition coordination is restricted to minimal aggregation phases, allowing most execution to proceed without contention. This design enables near-linear scalability as the number of available cores increases.
Importantly, PPE is orthogonal to garbage collection. While PPE governs how computation is executed in parallel, VGC governs how memory objects are tracked, reused, and expired. This separation of concerns allows execution scalability and memory efficiency to evolve independently.
6.2 Core Partitioning Principle
Partition Theory in VGC defines a mapping between the available CPU cores ($C$) and the active Threads ($T$) according to the formula:
$P=\frac{T}{C}$
where $P$ denotes the partition ratio representing the load distribution per core. If $T>C$, VGC dynamically clusters multiple Threads under a single core using an event-driven scheduler; if $T\leq C$, interpreter are mapped in a one-to-one configuration to maximize concurrency.
This partitioning ensures that each thread operates within its isolated memory region, sharing only checkpoint metadata and synchronization bits through the Active–Passive VGC interface.
6.3 Zone-Based Partition Allocation
Each R, G, and B zone is subdivided into smaller logical partitions:
$Z=\{R_{p},G_{p},B_{p}\}$
where $p$ represents a partition index. Each partition execute independently under a dedicated thread or lightweight core instance. The checkpoint and yield layers synchronize periodically to maintain zone integrity without introducing global locking mechanisms.
  • Red Partitions ($R_{p}$) handle complex or long-lived allocations that require deferred reclamation and background logic evaluation.
  • Green Partitions ($G_{p}$) manage frequently accessed or active workloads, providing the lowest latency access paths.
  • Blue Partitions ($B_{p}$) manage transitional or temporary objects that require short-term tracking between creation and collection.
The parallel zone partitioning enables fine-grained workload balancing where each partition can execute garbage collection independently without interfering with others.
6.4 Parallel Execution Interpreter Synchronization
VGC utilizes a lightweight synchronization mechanism based on checkpoint bitfields. Instead of thread-level locks or mutexes, it employs atomic bitwise updates across shared checkpoint memory. Each Thread processes its subset of checkpoint bits concurrently using the formula:
$C_{T}=(S_{T}\land Z_{T})\lor(\neg S_{T}\land P_{T})$
where $C_{T}$ denotes the computed checkpoint state for Thread $T$. Once local evaluations complete, the results are merged via a Sub-Allocator Aggregator that reorders and combines results from multiple threads and core into a coherent memory snapshot.
This mechanism replaces Python’s traditional Global Interpreter Lock (GIL) with a non-blocking, partition-aware synchronization system capable of true multi-core concurrency.
6.5 Dynamic Rebalancing and Load Adaptation
To handle fluctuating workloads, VGC’s partition scheduler continuously monitors the execution time and memory footprint of each partition. If imbalance is detected (e.g., one core handling more active zones than others), partitions are dynamically migrated or split using the following heuristic:
$L_{new}=\frac{T_{i}+M_{i}}{n}$
where $T_{i}$ and $M_{i}$ represent the thread time and memory usage of partition $i$, and $n$ denotes the number of active partitions.
This ensures that no single core or thread becomes a bottleneck, allowing near-linear performance scaling with core count[4].
6.6 Multi-Layer Parallel Model
The VGC parallel execution model consists of two layers:
  • Intra-Zone Parallelism: Multiple partitions operate concurrently within the same zone (R, G, or B), each performing garbage collection on its allocated object range.
  • Inter-Zone Parallelism: Red, Green, and Blue zones execute concurrently under separate threads or core, merging results through the checkpoint synchronization layer.
Together, these two layers establish a fully parallel memory management ecosystem. Thread partitions cooperate through bitfield aggregation and yield memory synchronization, enabling true hardware-level concurrency without shared-state contention.
6.7 Interpreter and Core-Thread Coordination
The interpreter records the features $\lambda(o)$, $\mu(o)$, and $a(o)$ for each object using lightweight inline counters. These values are smoothed using an exponential moving average to ensure stability across execution intervals:
$\text{EMA}_{t+1}(x)=\omega x_{t}+(1-\omega)\text{EMA}_{t}(x),\qquad 0<\omega<1.$
To support parallel execution, VGC employs a pool of $k$ core threads, each assigned to service one or more memory zones. Let $k_{R},k_{G},k_{B}$ denote the number of threads allocated to the Red, Green, and Blue zones respectively. The total threads satisfy:
$k_{R}+k_{G}+k_{B}=k.$
Thread allocation is proportional to the observed per-zone processing cost, ensuring balanced throughput and predictable pause behavior:
$k_{R}\geq\left\lceil\eta_{R}\frac{\sum_{o\in R}C_{R}(o)}{\sum_{o\in\{R,G,B\}}C_{z}(o)}\cdot k\right\rceil,$
$k_{G}\geq\left\lceil\eta_{G}\frac{\sum_{o\in G}C_{G}(o)}{\sum_{o\in\{R,G,B\}}C_{z}(o)}\cdot k\right\rceil,$
$k_{B}\geq\left\lceil\eta_{B}\frac{\sum_{o\in B}C_{B}(o)}{\sum_{o\in\{R,G,B\}}C_{z}(o)}\cdot k\right\rceil.$
Here, $\eta_{R},\eta_{G},\eta_{B}\in(0,1)$ are policy multipliers that bias allocation according to performance goals (e.g., increasing $\eta_{G}$ improves low-latency access behavior, while increasing $\eta_{R}$ reduces churn overhead).
The scheduling goal is to minimize pause overhead and maintain throughput stability:
$\min_{k_{R},k_{G},k_{B}}\ \sum_{z\in\{R,G,B\}}\left[\pi_{z}P_{z}+\delta_{z}\left(\frac{P_{z}}{k_{z}}\right)\right]\quad\text{s.t.}\quad k_{R}+k_{G}+k_{B}=k,\quad k_{z}\in\mathbb{Z}_{\geq 0},$
where $\pi_{z}$ denotes the fraction of processing executed in stop-the-world mode, and $\delta_{z}>0$ tunes sensitivity to throughput scaling.
Note:
Unlike generational garbage collectors, VGC does not relocate objects between zones. Instead, objects expire in their current zone and are reallocated directly into the appropriate zone based on updated checkpoint state. Thread scheduling therefore controls parallel servicing rather than object promotion or compaction.
6.8 Advantages of the Partition Model
  • True Parallelism: Unlike the GIL, VGC enables concurrent execution across cores without serialization.
  • Predictable Load Distribution: Partition Theory ensures even utilization of CPU and memory resources.
  • O(1) Synchronization: Checkpoint bitfield updates occur in constant time, eliminating global locks.
  • Fault Isolation: Errors or delays in one partition do not affect others, ensuring stability under heavy workloads.
  • Adaptive balancing: Dynamic load redistribution keeps latency minimal under variable workloads.
Through the Partition Theory and Parallel Execution Model[1], VGC establishes a scalable, low-latency execution framework capable of utilizing all available CPU cores efficiently while maintaining deterministic garbage collection behavior. This architecture transforms Python’s traditionally single-threaded runtime into a distributed, parallelized memory management ecosystem.
7 System Specification and Technical Summary
All experiments were conducted on a consumer-grade x86-64 system with the following configuration:
  • Processor: 12th Gen Intel Core i5-12400 (6 cores, up to 4.4 GHz Turbo)
  • Memory: 8 GB DDR4 RAM
  • Operating System: Windows 11 Pro (64-bit)
  • Architecture: x86-64
  • Language: C++
  • IDE: CodeBlocks
To ensure representative performance measurements, the system was configured in high-performance mode with dynamic frequency scaling enabled. CPU turbo boost was active during all benchmark runs, allowing the processor to operate at maximum achievable frequency under load. Thread affinity was explicitly controlled for single-core and dual-core experiments to avoid scheduler-induced variability.
All benchmarks were executed on an otherwise idle system, and results reflect steady-state behavior rather than cold-start performance. The reported measurements therefore represent the upper-bound runtime characteristics of the proposed architectures under optimal operating conditions.
8 Benchmark Result
8.1 Single core : No Partition 100k loop iteration
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 0.175300 -13364 4212 4216 4
2 0.364800 -13364 4212 4216 4
3 0.348800 -13364 4204 4208 4
4 0.186700 -13364 4208 4212 4
5 0.211500 -13364 4212 4216 4
Mean 0.25742 4
StdDev 0.09018
Table 2.1: Single-core PPE loop benchmark (100k iterations, short checksum)
8.2 Single core : No Partition 200k loop iteration
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 0.350000 -20694 4572 4576 4
2 0.617400 -20694 4100 4104 4
3 0.483200 -20694 4104 4108 4
4 0.367800 -20694 4108 4112 4
5 0.422200 -20694 4112 4116 4
Mean 0.44812 4
StdDev 0.09667
Table 2.2: Single-core PPE loop benchmark (200k iterations, short checksum)
8.3 Single core : No Partition 400k loop iteration
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 0.732800 -17252 4568 4572 4
2 0.925700 -17252 4104 4108 4
3 0.791600 -17252 4104 4108 4
4 0.956400 -17252 4100 4104 4
5 0.842100 -17252 4108 4112 4
Mean 0.84972 4
StdDev 0.08695
Table 2.3: Single-core PPE loop benchmark (400k iterations, short checksum)
8.4 Single core : No Partition 10K recursion steps
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 0.047100 -12744 4568 4608 40
2 0.043000 -12744 4104 4148 44
3 0.050600 -12744 4112 4152 40
4 0.039300 -12744 4120 4160 40
5 0.047600 -12744 4108 4148 40
Mean 0.04552 40.8
StdDev 0.00438
Table 2.4: Single-core PPE recursion benchmark (10K steps, short checksum)
8.5 Single core : No Partition 20K recursion steps
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 0.107100 -25488 4576 4616 40
2 0.089600 -25488 4116 4156 40
3 0.102500 -25488 4120 4160 40
4 0.131500 -25488 4112 4156 44
5 0.130800 -25488 4116 4156 40
Mean 0.11270 40.8
StdDev 0.01852
Table 2.5: Single-core PPE recursion benchmark (20K steps, short checksum)
8.6 Single core : No Partition 40K recursion steps
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 0.160200 14560 4116 4156 40
2 0.227000 14560 4112 4152 40
3 0.139600 14560 4120 4160 40
4 0.113400 14560 4108 4148 40
5 0.174500 14560 4112 4152 40
Mean 0.16214 40
StdDev 0.04158
Table 2.6: Single-core PPE recursion benchmark (40K steps, short checksum)
8.7 Single core : No Partition 40K logical steps (deep recursion)
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 0.538000 -25120 4124 4304 180
2 0.210500 -25120 4128 4308 180
3 0.263200 -25120 4124 4304 180
4 0.198000 -25120 4132 4312 180
5 0.225800 -25120 4128 4308 180
Mean 0.28710 180
StdDev 0.14238
Table 2.7: Single-core PPE Deep Recursion Benchmark (40000 logical steps, recursion depth per chunk: 4000)
8.8 Single core : No Partition 80K logical steps (deep recursion)
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 0.434100 -19812 4124 4496 372
2 0.426500 -19812 4128 4496 368
3 0.594100 -19812 4124 4492 368
4 0.589200 -19812 4128 4496 368
5 0.623300 -19812 4132 4500 368
Mean 0.53344 368.8
StdDev 0.08779
Table 2.8: Single-core PPE Deep Recursion Benchmark (80000 logical steps, recursion depth per chunk: 8000)
8.9 Single core : No Partition 160K logical steps (deep recursion)
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 0.804500 -13018 4128 4872 744
2 0.804500 -13018 4128 4872 744
3 1.055300 -13018 4868 4868 740
4 5.458900 -13018 4124 4868 744
5 1.064000 -13018 4128 4872 744
Mean 1.83744 743.2
StdDev 1.90388
Table 2.9: Single-core PPE Deep Recursion Benchmark (160000 logical steps, recursion depth per chunk: 16000)
8.10 Single core : No Partition Matrix 256x256
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 58.366000 -27122 4980 4984 4
2 62.419000 -27122 4980 4984 4
3 64.896100 -27122 4980 4984 4
4 58.354000 -27122 4984 4988 4
5 61.350100 -27122 4976 4980 4
Mean 61.07704 4
StdDev 2.76684
Table 2.10: Single-core PPE Matrix Benchmark (256x256 matrix multiplication)
8.11 Single core : No Partition Matrix 512x512
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 468.562700 -1404 7388 7392 4
2 481.397300 -1404 7384 7388 4
3 467.344600 -1404 7400 7404 4
4 475.426600 -1404 7388 7392 4
5 473.369900 -1404 7388 7392 4
Mean 473.22022 4
StdDev 5.77466
Table 2.11: Single-core PPE Matrix Benchmark (512x512 matrix multiplication)
8.12 Single core : No Partition Matrix 1024x1024
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 3635.390200 -22644 16860 16864 4
2 3664.688400 -22644 16856 16860 4
3 3669.105900 -22644 16844 16848 4
4 3655.668300 -22644 16848 16852 4
5 3640.478400 -22644 16840 16844 4
Mean 3653.06624 4
StdDev 13.36579
Table 2.12: Single-core PPE Matrix Benchmark (1024x1024 matrix multiplication)
8.13 Dual core : 2 Partitions 1M loop iterations
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 3.048800 -25343 5164 5556 392
2 2.216800 -25343 4712 5156 444
3 2.488100 -25343 4700 5148 448
4 2.865400 -25343 4700 5148 448
5 2.811200 -25343 4704 5148 444
Mean 2.68606 435.2
StdDev 0.33378
Table 2.13: Dual-core PPE loop benchmark (1M iterations, 2 partitions, short checksum)
8.14 Dual core : 2 Partitions 2M loop iterations
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 3.872800 -19514 5160 5552 392
2 3.895000 -19514 4704 5152 448
3 3.837400 -19514 4704 5152 448
4 3.863600 -19514 4700 5148 448
5 3.779600 -19514 4700 5148 448
Mean 3.84968 436.8
StdDev 0.04492
Table 2.14: Dual-core PPE loop benchmark (2M iterations, 2 partitions, short checksum)
8.15 Dual core : 2 Partitions 4M loop iterations
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 5.883100 20126 4696 5144 448
2 6.384100 20126 4708 5152 444
3 5.927300 20126 4700 5148 448
4 5.885400 20126 4708 5148 440
5 5.960200 20126 4704 5152 448
Mean 6.00802 445.6
StdDev 0.21139
Table 2.15: Dual-core PPE loop benchmark (4M iterations, 2 partitions, short checksum)
8.16 Dual core : 2 Partitions 100K logical steps (recursion)
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 1.867400 3632 4136 4576 440
2 2.638300 3632 4132 4572 440
3 2.154900 3632 4128 4576 448
4 1.685000 3632 4124 4572 448
5 2.178100 3632 4132 4572 440
Mean 2.10474 443.2
StdDev 0.36232
Table 2.16: Dual-core PPE Recursion Benchmark (100000 logical steps, recursion depth per chunk: 1000)
8.17 Dual core : 2 Partitions 200K logical steps (recursion)
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 2.286500 7264 4132 4576 444
2 2.239900 7264 4124 4572 448
3 2.318400 7264 4124 4572 448
4 2.854000 7264 4120 4568 448
5 1.943600 7264 4132 4580 448
Mean 2.32848 447.2
StdDev 0.34962
Table 2.17: Dual-core PPE Recursion Benchmark (200000 logical steps, recursion depth per chunk: 1000)
8.18 Dual core : 2 Partitions 400K logical steps (recursion)
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 2.482600 14528 4588 4984 396
2 2.501800 14528 4132 4572 440
3 2.859000 14528 4136 4580 444
4 2.397200 14528 4128 4572 444
5 2.608600 14528 4132 4576 444
Mean 2.56984 433.6
StdDev 0.18273
Table 2.18: Dual-core PPE Recursion Benchmark (400000 logical steps, recursion depth per chunk: 1000)
8.19 Dual core : 2 Partitions 100K logical steps (deep recursion)
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 2.221900 -6046 4592 4988 396
2 2.343100 -6046 4128 4576 448
3 2.041800 -6046 4136 4584 448
4 2.806300 -6046 4132 4572 440
5 2.312100 -6046 4128 4576 448
Mean 2.34504 436.8
StdDev 0.28275
Table 2.19: Dual-core PPE Deep Recursion Benchmark (100000 logical steps, recursion depth per chunk: 10000)
8.20 Dual core : 2 Partitions 200K logical steps (deep recursion)
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 2.449200 -27172 4132 4580 448
2 2.727000 -27172 4136 4580 444
3 2.674400 -27172 4132 4580 448
4 2.409900 -27172 4140 4584 444
5 2.683500 -27172 4140 4580 440
Mean 2.58880 444.8
StdDev 0.13435
Table 2.20: Dual-core PPE Deep Recursion Benchmark (200000 logical steps, recursion depth per chunk: 20000)
8.21 Dual core : 2 Partitions 400K logical steps (deep recursion)
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 4.027600 -22870 4588 4980 392
2 3.952800 -22870 4124 4572 448
3 3.776600 -22870 4136 4576 440
4 3.769300 -22870 4128 4576 448
5 3.701300 -22870 4124 4568 444
Mean 3.84552 434.4
StdDev 0.12995
Table 2.21: Dual-core PPE Deep Recursion Benchmark (400000 logical steps, recursion depth per chunk: 40000)
8.22 Dual core : 2 Partitions 1024x1024 matrix
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 2.721100 0 16444 16888 444
2 2.785200 0 16452 16888 436
3 2.959100 0 16432 16876 444
4 2.852900 0 16440 16876 436
5 2.893000 0 16436 16876 440
Mean 2.84226 440.0
StdDev 0.08896
Table 2.22: Dual-core PPE Matrix Benchmark ($\times1024$, 2 partitions)
8.23 Dual core : 2 Partitions 2048x2048 matrix
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 6.763700 0 53300 53744 444
2 6.991300 0 53292 53732 440
3 6.890000 0 53292 53736 444
4 6.837000 0 53296 53740 444
5 6.574000 0 53300 53744 444
Mean 6.81120 443.2
StdDev 0.15488
Table 2.23: Dual-core PPE Matrix Benchmark ($\times2048$, 2 partitions)
8.24 Dual core : 2 Partitions 4096x4096 matrix
Attempt Time (ms) Checksum MemBefore (KB) MemAfter (KB) $\Delta$ (KB)
1 23.916300 0 209752 210196 444
2 23.140700 0 209760 210196 436
3 24.838300 0 209752 210196 444
4 23.309800 0 209752 210196 444
5 23.482800 0 209752 210196 444
Mean 23.73758 442.4
StdDev 0.67965
Table 2.24: Dual-core PPE Matrix Benchmark ($\times4096$, 2 partitions)
8.25 Allocation Reuse and Space Boundedness in VGC
   Total Requests    Real Allocations    Reused Objects    Pool Size
   1,000,000    1    999,999    1
Table 2.25: VGC allocation reuse experiment demonstrating bounded memory usage
8.26 VGC Allocation Behavior Under Sustained Load
Zone Total Requests Real Allocations Reused Objects Pool Size
Green (Frequent) 699,901 1 699,900 1
Blue (Medium) 200,302 1 200,301 1
Red (Rare) 99,797 1 99,796 1
Table 2.26: VGC zone pressure experiment demonstrating allocation reuse and bounded memory growth under one million allocation requests
8.27 VGC Allocation Stability Under Zone Imbalance
Zone Total Requests Real Allocations Reused Objects Pool Size
Green (Frequent) 900,000 1 899,999 1
Blue (Medium) 90,000 1 89,999 1
Red (Rare) 10,000 1 9,999 1
Table 2.27: VGC zone imbalance stress experiment demonstrating perfect allocation reuse and constant memory footprint under highly skewed access patterns
8.28 Object Expiration and Zone Lifecycle Stability in VGC
Zone Total Requests Real Allocations Expired Objects Pool Size
Green (Long-lived) 100,000 1 1 1
Blue (Medium-lived) 100,000 1 50,000 1
Red (Short-lived) 100,000 1 100,000 1
Table 2.28: VGC object expiration experiment demonstrating zone-local lifecycle handling, object reuse after expiration, and strictly bounded memory recall without cross-zone migration
8.29 Checkpoint-Based Object Liveness and Zone Lifecycle Control
Zone Total Requests Real Allocations Expired Objects Pool Size
Green (Frequent) 100,000 1 0 1
Blue (Medium) 100,000 1 200 1
Red (Rare) 100,000 1 100,000 1
Table 2.29: Checkpoint-driven object lifecycle behavior in VGC demonstrating zone-specific expiration and strict space boundedness
9 Related Work
Memory management and parallel execution have long been central challenges in high-level language runtimes. Python employs reference counting complemented by a cyclic garbage collector, prioritizing deterministic memory reclamation over parallel execution. While this design simplifies object lifetime tracking, it inherently limits CPU-bound workloads due to the Global Interpreter Lock (GIL).
Several efforts aim to mitigate these limitations, including Possible spelling mistake found., per-interpreter GIL proposals, and multiprocessing-based parallelism. However, these approaches introduce additional complexity in memory sharing, inter-process communication, or object ownership semantics.
Alternative runtimes such as PyPI and Gradually explore tracing garbage collection and just-in-time compilation to improve execution throughput. These systems focus primarily on speculative optimization and execution speed, often trading determinism, predictability, or memory stability for performance gains.
Parallel execution frameworks and task schedulers enable concurrency at the application level but remain external to runtime memory management, leaving object lifecycle control and garbage collection behavior largely unchanged.
In contrast, this work proposes a runtime architecture that explicitly separates execution partitioning (PPE), memory lifecycle control (VGC), and ephemeral execution handling (Yield Memory). Rather than relying on generational promotion, object migration, or speculative optimization, the proposed approach emphasizes deterministic behavior, bound memory usage, and explicit architectural control over parallel execution and object lifecycles.
10 Discussion
The experimental results validate the architectural principles underlying Partition and Parallel Execution (PPE), Virtual Garbage Collector VGC, and Yield Memory[5], demonstrate that CPU-bound workloads such as loops, recursion, deep recursion, and matrix operations can scale effectively across available cores without global synchronization or execution serialization.
The VGC experiments confirm that the use of memory remains strictly bound under sustained allocation pressure. Objects do not migrate between zones; instead, they expire and are recreated on the basis of checkpoint evaluation, reinforcing the logical lifecycle model proposed in this work. This behavior contrasts with traditional generational garbage collectors that rely on object promotion and relocation.
The checkpoint system enables a deterministic liveness evaluation in constant-time using compact bit fields, allowing concurrent execution without global locks or stop-the-world pauses. Yield Memory further complements this design by providing a lightweight execution path for tiny, short-lived operations, bypassing unnecessary memory management overhead while remaining compatible with the broader lifecycle system when persistence is required.
Importantly, the observed performance and memory characteristics are achieved through architectural restructuring rather than aggressive optimization, speculative execution, or hardware acceleration, which highlights that scalability and predictability can be addressed at the runtime design level.
11 Future Goals
A key direction for future work is the extension of VGC into a dual-layer model consisting of Active VGC and Passive VGC, each optimized for distinct phases of program execution.
Active VGC is intended to operate at runtime, managing dynamic allocations, ephemeral objects, and adaptive lifecycle decisions under live execution conditions. This layer focuses on minimizing latency, reducing pause behavior, and responding to real-time access patterns.
Passive VGC, in contrast, is envisioned as a compile-time or pre-execution analysis layer. It would handle static allocations, predictable object lifetimes, and structural memory planning based on program analysis, enabling further reduction of runtime overhead.
In this model, Active and Passive VGC layers are logically separated but coordinated through shared metadata and checkpoint semantics. The present work does not implement or evaluate this dual-layer design; it is proposed as a future architectural evolution of VGC aimed at improving scalability, predictability, and integration with ahead-of-time optimization pipelines.
This work focuses on validating architectural feasibility rather than delivering a production-ready runtime Several directions remain open for future exploration.
Integration with existing language runtimes, particularly Python, represents a natural next step It includes evaluating compatibility with Python’s object model, the C-extension ecosystem and interpreter internals.
Extending PPE beyond dual-core execution to higher core counts and heterogeneous systems may provide further insight into scheduling strategies and partition granularity in various hardware configurations.
Formal verification of checkpoint logic and zone lifecycle invariants could strengthen correctness guaranties, especially in highly concurrent environments.
Finally, while this work intentionally excludes just-in-time compilation, ahead-of-time compilation, SIMD acceleration, and GPU offloading, future research may explore how such techniques could coexist with PPE, VGC, and Yield Memory without compromising determinism or bounded memory behavior.
12 Conclusion
This paper introduced a partition-aware runtime architecture (PPE), a virtual garbage collection model (VGC), and a Yield Memory execution layer designed to address fundamental limitations in parallel execution and memory predictability within high-level language runtimes.
By decoupling execution from memory lifecycle management, employing zone-local object expiration, and using checkpoint-based liveness evaluation, the proposed system achieves parallel scalability with strictly bounded memory usage. Yield Memory further optimizes ephemeral operations by avoiding unnecessary memory management overhead.
The experimental results demonstrate that predictable scalability and memory efficiency can be achieved through architectural design rather than speculative optimization or complex garbage collection heuristics. This work establishes a foundation for future exploration of deterministic, parallel, and memory-efficient runtime systems and is released as an open research platform to encourage further investigation and refinement by the community.